OpenStudy (anonymous):

1. Write the equation of the line with a slope of 2/3 and a y-intercept of 5 in general form. Clearly show and explain each step

OpenStudy (anonymous):

well you start with slope-intercept form and you know what the slope and the y-intercept are so you write Y=2/3x+5 To change it to General form just move all of that to look like Ax+By+C=0 Ax=2/3x By=Y C=5 so then it looks like this 2/3x+y+5=0